OpenStudy (anonymous):

A curve with a y-intercept of 4 and passing through the point (2,e^2) has the property that the slope of the curve at any point (x,y) is directly proportional to the y-coordinate of the point. What is an equation of this curve?

OpenStudy (anonymous):

The slope \(y'\) is directly proportional to the value of \(y\) at any given point, which means \[y'=ky\] for some real \(k\). This looks like the setup for solving a differential equation, so I'm going to operate under that assumption. You have two initial conditions (you'll need both). \[y'=ky~~\iff~~\frac{dy}{y}=k\,dx~~\iff~~\int\frac{dy}{y}=k\int dx\] Integrating and solving for \(y\) yields an exponential curve. You can find the value of the constant of integration by using the first condition, and the value of \(k\) by using the second.

OpenStudy (anonymous):

so \[y=e ^{kx+C}\] separate that to \[y=e ^{kx}+e ^{C}\] let \[e ^{C}=A\] \[y=Ae ^{kx}\] plug in the point (0,4) to find \[A=4\] then plug in the point (2,e^2) to find \[k=1-\ln 2\] and the final equation should be \[y=4e ^{(1-\ln 2)x}\] Is that right?

OpenStudy (anonymous):

Your separation had a slight mistake, but you corrected it: \(y=e^{kx+C}=e^Ce^{kx}=Ae^{kx}\). Aside from that, yes, that's right!

OpenStudy (anonymous):

Thank you!!